OpenStudy (anonymous):

The average tuition price of two colleges in Georgia is $7,700. In New York, the average tuition price of three colleges is $9,400. What is the weighted mean tuition price for all colleges together? Please show work for calculating the weighted mean. Weighted mean _

OpenStudy (anonymous):

what does weighted mean mean?

OpenStudy (anonymous):

if the average of two colleges in georgia is 7700, how much does it cost to attend both schools total?

OpenStudy (anonymous):

nooooo idea

OpenStudy (anonymous):

multiply the number of schools by average cost

OpenStudy (anonymous):

for both states

OpenStudy (anonymous):

then add it up and divide it by the amount of schools there are- which is 5

OpenStudy (anonymous):

you can think of it this way each of the two schools in georgia cost 7,700 and each of the three schools in ny cost 9,400 so the total of all five schools is \[2\times 7,700+3\times 9,400\] and the average is \[\frac{2\times 7,700+3\times 9,400}{5}\]

OpenStudy (anonymous):

thank you.:)

OpenStudy (anonymous):

thats a lot simplier...wow

OpenStudy (anonymous):

you must study math all the time

OpenStudy (anonymous):

id prob rather shoot myself

OpenStudy (anonymous):

it must be over soon i imagine

OpenStudy (anonymous):

one last section and then im done for good.

OpenStudy (anonymous):

a lot fo this feel irrelevant, though

OpenStudy (anonymous):

done for good? or for the moment?

OpenStudy (anonymous):

It creates unneeded frustration

OpenStudy (anonymous):

yeah, that is always true about math but then it comes back and you think "oh, that is what it was about..."

OpenStudy (anonymous):

Like you dont know how many people get to the point of crying just because they don't understand this and since math is such a big deal, people are taught that basically if you dont have the skills in math, you're screw**

OpenStudy (anonymous):

*done for good

OpenStudy (anonymous):

lol yeah it can be stressful but you have to admit that last problem was not such a big deal, was it?

OpenStudy (anonymous):

not really, but these other ones.... i'm not good at charts

OpenStudy (anonymous):

just think, when you are done, you can come back here and answer questions yourself

OpenStudy (anonymous):

it is kind of important to be able to read a graph right?

OpenStudy (anonymous):

maybe:)

OpenStudy (anonymous):

of course, but it can be difficult

OpenStudy (anonymous):

you will be back helping others i am sure

OpenStudy (anonymous):

math makes me feel sad because I'm right-brained

OpenStudy (anonymous):

i can see that from your picture but it should not make you sad, just like right brained stuff does not make me say, even if i don't understand it some times